OpenStudy (anonymous):

Let f(x) be the square of the distance from the point (2,1) to a point (x,3x+2) on the line y=3x+2. Find the minimum value of f(x) using either algebraic techniques or graphing techniques.

OpenStudy (freckles):

Did you find the distance square between the points that were given?

OpenStudy (anonymous):

sqrt((2-x)^2+(1-3x+2)^2)

OpenStudy (freckles):

well if you wanted to write that way the only bad thing I see is you didn't write 1-3x-2 in that one ( )

OpenStudy (freckles):

in order words you would have \[d=\sqrt{(2-x)^2+(1-[3x+2])^2}\] you need to distribute that -

OpenStudy (freckles):

\[d=\sqrt{(2-x)^2+(1-3x-2)^2}\]

OpenStudy (freckles):

Anyways distance squared would look like this: \[d^2=(2-x)^2+(1-3x-2)^2\] And it told us to rename this f(x) so \[f(x)=(2-x)^2+(1-3x-2)^2\]

OpenStudy (freckles):

Now you should realize the degree here is 2 so you should be able to write it in y=ax^2+bx+c or y=a(x-h)^2+k <---which will give the min of the parabola

OpenStudy (freckles):

You need simplify 1-3x-2 first

OpenStudy (freckles):

then do some multiply and combining of like terms

OpenStudy (triciaal):